# Dissipative spatiotemporal soliton in a driven waveguide laser

### v1

@@ 2.2 Variational approximation
The study of STDS of Eq. (2) is based on analytical and numerical approaches. The first uses the well-known variational approximation (VA) [55]. The generating Lagrangian for a non-dissipative part of (2) is
$\mathscr{L}=\frac{i}{2}\left(\psi^{*}\partial_{Z}\psi-\psi\partial_{Z}\psi^{*}\right)+\frac{1}{2}\left(\left|\partial_{X}\psi\right|^{2}+\left|\partial_{Y}\psi\right|^{2}+\left|\partial_{t^{\prime}}\psi\right|^{2}\right)-$
$\left|\psi\right|^{4}+\left(X^{2}+Y^{2}\right)\left|\psi\right|^{2}.$ (3)
The dissipative factors can be described in the form of a “force” source term
@@
in the Euler-Lagrange equation
@@
in agreement with Kantorovitvh’s method (see [55, 15, 53] and references herein).
It is convenient to assume the axial symmetry, which corresponds to the symmetry of a cylindrical waveguide, and transit to a radial coordinate $R=\sqrt{X^{2}+Y^{2}}$. The next assumption is a Gaussian-sech approximation for an STDS profile:
@@
where the $Z-$dependent STDS parameters of f in Eq. (5) are: $\alpha$ amplitude, $\phi$ phase, $\chi$ and $\theta$ are the temporal and spatial chirps, respectively; $\rho$ and $\Upsilon$ are the beam size and STDS duration, respectively. After some algebraic manipulations, one can obtain the expressions for the steady-state (i.e., $Z-$independent) STDS parameters (the phase $\phi$ is irrelevant in our context):
$\rho^{2}=\frac{1}{384\pi^{2}\kappa\tau\Upsilon^{2}}\times$ (7)
$\biggl{[}15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}-15\right)-$
$64\tau\left(15\tau+2\pi^{2}\left(\tau+3\Lambda^{\prime}\Upsilon^{2}\right)\right)\biggr{]},$
$\chi=\frac{\sqrt{\Upsilon^{4}\left(64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225\right)}-15\Upsilon^{4}}{16\pi^{2}\tau\Upsilon^{4}},$ (8)
$\alpha_{0}^{2}=\frac{-6\kappa^{2}\rho^{8}-6\rho^{4}+6}{2\rho^{2}},$
$\theta=-\frac{\kappa\rho^{2}}{2}.$
@@
$144\left(\pi^{2}-25\right)\tau^{2}+\left(16\left(3+\pi^{2}\right)\tau^{2}+45\right)\times$ (9)
$\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}+96\pi^{4}\Upsilon^{4}\omega+$
$+\frac{2^{13}3^{3}\pi^{4}\kappa\tau^{3}\Upsilon^{4}\left(-\frac{A^{4}}{3\times 2^{15}\pi^{8}\kappa^{2}\tau^{4}\Upsilon^{8}}-\frac{3A^{2}}{2\pi^{4}\kappa^{2}\tau^{2}\Upsilon^{4}}+1\right)}{A}=675,$ (10)
$A=64\tau\left(2\pi^{2}\left(3\Lambda^{\prime}\Upsilon^{2}+\tau\right)+15\tau\right)-$
$15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon^{4}\omega\right)+225}-15\right)$
must be solved numerically.
@@ 2.3 STDS parameters
Figs. 2-5 show the STDS width $\Upsilon$ and peak intensity $\alpha_{0}^{2}$ obtained from Eqs. (7-9). $\Upsilon$ grows, and $\alpha_{0}^{2}$ decreases with $\kappa$ when a spectral dissipation $\tau$ is sufficiently low (Figs. 2, 3). Such a tendency corresponds to that of a DKLM regime with $\omega=0$ and manifests the growth of spatial mode loss with an aperture squeezing [55].
In the vicinity of a “threshold” ($\Lambda^{\prime}\to$0), the additional longitudinal (3D) trapping defined by $\omega$ noticeably squeezes an STDS with the simultaneous peak power growth (Fig. 3). The additional solution branch in the region of small $\kappa$ (Fig. 2) disappears with $\Lambda^{\prime}\to$0. In general, the STDS intensity growth and its width shortening with $\omega$ could result from an appearance of the external confinement along the $t-$ coordinate.
The spectral dissipation increases with $\tau$. That leads to an STDS collapse for $\kappa\to|\Lambda^{\prime}|$ when the net-loss coefficient $\Lambda^{\prime}$ is more significant in comparison with Fig. 3. Such a collapse is asymmetrical in space and time. That is, STDS spreads along $t$ and squeezes on $R$. A longitudinal mode selection growth with $\tau$ (i.e., STDS spreading in the time domain) parallels a transverse mode selection (STDS spatial squeezing) in the spatial domain.
The solutions bifurcate close to the threshold $\Lambda^{\prime}$ (Fig. 5). Such behavior leads to the thought of stability loss here. These oddities require numerical simulations of STDS dynamics presented below.
Figure 2: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=$-0.01, $\tau=$0.01.
Figure 3: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=-$0.001, $\tau=$0.01.
Figure 4: Dimensionless pulse width $\Upsilon$ (left axis) and intensity $\alpha_{0}^{2}$ (right axis) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=-$0.01, $\tau=$0.1.
@@
Figure 5: Dimensionless pulse width $\Upsilon$ (a) and intensity $\alpha_{0}^{2}$ (b) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=-$0.001, $\tau=$0.1.
2.4 STDS stability
@@
We performed large-scale numerical simulations of Eq. (2) based on the COMSOL Multiphysics software to analyze the stability of the analytical solutions based on the VA. The characteristic case is illustrated in Fig. 6, where the numerical STDS profile corresponds to one of the VA solutions from Fig. 4. The red line shows the analytical and numerical solutions shown by the black curve. One can see an oscillatory approach from the last to the first, analogous to the damping of excitations in a weakly dissipative BEC [69]. Interestingly, the peak intensity of the numerical solution is slightly lower than the analytical one. The difference in the exact solution from the Gaussian ansatz for VA could explain it.
[…]
Besides the stable STDS solutions, the numerical simulations demonstrate three main destabilization scenarios. The first is the STDS decay with the $|\Lambda^{\prime}|$-decrease. The second is the collapse-like behavior for small $\kappa$, $\omega$, and $\tau$. Both demonstrate the importance of dissipative factors (i.e., dissipative confinement and spectral filtering) for STDS stability.
@@
The third scenario is the DS splitting with possible oscillations between multiple pulse states. Fig. 7 shows such regimes. The splitting in a local time domain could be explained by weaker confinement in the time domain due to lower $\omega$. However, a more precise analysis of the STDS requires considering the gain saturation dynamics.
Figure 7: Contour plot of the STDS peak intensity $\alpha_{0}^{2}$ at $Z=$40000 (a) and 35000 (b) for $\omega=10^{-5}$ (a) and $2\times 10^{-6}$ (b), $\kappa=$0.005, $\Lambda^{\prime}=$-0.01, and $\tau=$0.1.
@@ 2.5 STDS stability under the saturable gain dynamics
To take into account the dynamical gain saturation, it is convenient to use the Frantz-Nodvik equation [70]:
@@
where $G_{ss}$ is a small signal gain, $T_{r}$ is a gain relaxation time, $E_{s}=h\nu/S\sigma_{em}$ is a gain saturation energy ($\nu$ is a central gain-band wavelength coinciding with the STDS carrier frequency, $S$ is a beam area, $\sigma_{em}$ is a stimulated emission cross-section). Specifically, we base on the characteristics of the Cr:ZnS waveguide laser [71]. Table 3 presents the corresponding dimensional laser parameters.
Table 3: Correspondences between photonics and BEC
Cr:ZnS laser parameters (see Table 1) values
gain relaxation time $T_{r}$, $\mu s$ 4.9
@@
central wavelength $\lambda$, $\mu$m 2.3
@@
$n_{0}$ 2.2629
$\delta$ 0.0077
numerical aperture $NA=\sqrt{n_{1}^{2}-n_{0}^{2}}$ $\approx 0.19$
@@
The calculations demonstrate a visible effect of the gain saturation on the DS dynamics. For a small level of linear loss $\Lambda$, some characteristic profiles of STDS are shown in Fig. 8. As one can see, a sufficiently large $\omega$ stabilizes DS ($a$), but the driving frequency decrease leads to the STDS shape distortion, which is periodically asymmetrical along $t$-coordinate. Oscillations accompany these distortions and result in STDS decay. STDS cannot be stabilized for the parameters considered for a smaller $\Lambda$ (Fig. 5).
@@
The region of STDS stability is shown in Fig. 9 for a larger linear loss coefficient $\Lambda$. The STDS decay confines this region for larger $\kappa$, so there are no stable pulses for $\kappa\gtrsim 0.006$. The difference with the analytical results lies in the reduction of a net gain required for the soliton formation. Such a gain reduction is a direct consequence of the gain saturation. For smaller $\kappa$, the main destabilization factor is periodical pulse splitting, which alternates with a single-pulse regime (Figs. 8 (b-d) and inset in Fig. 9). The asymmetry of the STDS complex appears periodically due to dynamic gain saturation. When $\kappa$ approaches zero, very strong STDS power and width oscillations appear. They reach more than two orders of magnitude, so STDS is periodically “reborn”. This phenomenon confines the stability region from below on $\kappa$.
Figure 9: Region of STDS stability on ($\kappa$$\omega$)-plane under the action of dynamical gain saturation. $\Lambda=$0.1, and other parameters as in Fig. 8. Inset shows the $|\alpha_{0}|^{2}$ contour plot for $\kappa=5\times 10^{-4}$, $\omega=\times 10^{-6}$.

### v2

@@ 2.2 Variational approximation
The study of STDS of Eq. (2) is based on analytical and numerical approaches. The first uses the well-known variational approximation (VA) Karlsson et al. (1992); Yu et al. (1995); Malomed (2002); Kalashnikov and Wabnitz (2020); Parra-Rivas et al. (2023). The generating Lagrangian density for a non-dissipative part of (2) is
$\mathscr{L}=\frac{1}{2}\left\{i\left(\psi^{*}\partial_{Z}\psi-\psi\partial_{Z}\psi^{*}\right)+\left(\left|\partial_{X}\psi\right|^{2}+\left|\partial_{Y}\psi\right|^{2}+\left|\partial_{t^{\prime}}\psi\right|^{2}\right)\right\}-$
$\frac{1}{2}\left|\psi\right|^{4}+\frac{1}{2}\left(X^{2}+Y^{2}\right)\left|\psi\right|^{2}.$ (3)
The dissipative factors can be described in the form of the Rayleigh dissipative function
@@
in the Euler-Lagrange-Kantorovich equation
@@
in agreement with Kantorovich’s method, which is a generalization of the Rayleigh-Ritz method Kantorovich and Krylov (1958); Chávez Cerda et al. (1998). In Eq. (5), the variation is made over the $Z-$dependent parameters ${\mathop{\rm f}\nolimits}$ of some ansatz, which will be introduced below. The reduced Lagrangian
[…]
is evaluated after substituting an ansatz in $L$ and $Q$.
It is convenient to assume the axial symmetry, which corresponds to the symmetry of a cylindrical waveguide, and transit to a radial coordinate $R=\sqrt{X^{2}+Y^{2}}$. The next assumption is a Gaussian-sech ansatz for Eqs. (5,6) Yu et al. (1995):
@@
where the $Z-$dependent STDS parameters of f in Eq. (5) are: $\alpha$ amplitude, $\phi$ phase, $\chi$ and $\theta$ are the temporal and spatial chirps, respectively; $\rho$ and $\Upsilon$ are the beam size and STDS duration, respectively. We use this ansatz as an approximation for a pulse with the near-Gaussian transverse spatial profile, i.e., the lowest-order spatial mode in the parabolic potential of Eq. (3), and a sech-shaped pulse with the quadratic phase for a temporal profile. The last was widely used for VA. Other ansatzs with different temporal phase profiles were also considered (e.g., see Malomed (2002). In our case, the temporal shape of ansatz is conditioned by the following. This shape corresponds to a dissipative soliton solution $\operatorname{sech}(t^{\prime}/\Upsilon)^{1-i\psi}$ of Eq. (2) without a confining potential Maimistov and Skliarov (1987) in a limit when the exact phase profile $-\psi\ln{(\operatorname{sech}(t^{\prime}/\Upsilon)})\approx\psi t^{\prime 2}/2\Upsilon^{2}$ (here, $\psi=2\Upsilon^{2}\chi$ is a dimensionless chirp).
After some algebraic manipulations (see Appendix A and Kalashnikov (2024)), one can obtain the expressions for the steady-state (i.e., $Z-$independent) STDS parameters (there are two solutions for the chirp parameter $\chi$, but one of them is nonphysical):
$\rho_{0}^{2}=\frac{1}{384\pi^{2}\kappa\tau\Upsilon_{0}^{2}}\times$
$\biggl{[}15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon_{0}^{4}\omega\right)+225}-15\right)-$
$64\tau\left(15\tau+2\pi^{2}\left(\tau+3\Lambda^{\prime}\Upsilon_{0}^{2}\right)\right)\biggr{]},$
$\chi_{0}=\frac{\sqrt{\Upsilon_{0}^{4}\left(64\tau\left(30\tau-\pi^{4}\Upsilon_{0}^{4}\omega\right)+225\right)}-15\Upsilon_{0}^{4}}{16\pi^{2}\tau\Upsilon_{0}^{4}},$ (8)
$\alpha_{0}^{2}=\frac{-6\kappa^{2}\rho_{0}^{8}-6\rho_{0}^{4}+6}{2\rho^{2}},$
$\theta_{0}=-\frac{\kappa\rho_{0}^{2}}{2}.$
@@
$144\left(\pi^{2}-25\right)\tau^{2}+\left(16\left(3+\pi^{2}\right)\tau^{2}+45\right)\times$
$\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon_{0}^{4}\omega\right)+225}+96\pi^{4}\Upsilon_{0}^{4}\omega\tau+$
$+\frac{2^{13}3^{3}\pi^{4}\kappa\tau^{3}\Upsilon_{0}^{4}\left(-\frac{A^{4}}{3\times 2^{15}\pi^{8}\kappa^{2}\tau^{4}\Upsilon_{0}^{8}}-\frac{3A^{2}}{2\pi^{4}\kappa^{2}\tau^{2}\Upsilon_{0}^{4}}+1\right)}{A}=675,$ (9)
$A=64\tau\left(2\pi^{2}\left(3\Lambda^{\prime}\Upsilon_{0}^{2}+\tau\right)+15\tau\right)-$
$15\left(\sqrt{64\tau\left(30\tau-\pi^{4}\Upsilon_{0}^{4}\omega\right)+225}-15\right)$
must be solved numerically. Below, we will omit the 00-subscribe for steady-state (fixed-point) solutions characterizing STDS.
@@ 2.3 STDS parameters
Figs. 2-4 show the STDS width $\Upsilon$ and peak intensity $\alpha_{0}^{2}$ obtained from Eqs. (8, 9). $\Upsilon$ grows, and $\alpha_{0}^{2}$ decreases with $\kappa$. Such a tendency corresponds to that of a DKLM regime with $\omega=0$ and manifests the growth of spatial mode loss with a soft aperture squeezing $\kappa=G_{0}w_{p}^{-2}$ Kalashnikov and Wabnitz (2020). The nonzero $\omega$ contributes to this behavior and squeezes the $\kappa$-range where the STDS exists. It could be explained by an additional loss due to confinement in the $t$-domain. Insets in Fig. 4 demonstrate the pulse spatio-temporal and spectral profiles corresponding to the ansatz (7) (see Appendix B).
The solutions for $\omega\neq$0 bifurcate close to the threshold $\Lambda^{\prime}\to$0 (Figs. 3,4). The new solutions with the opposite sign of the chirp $\chi$ appear. The dependencies of $\Upsilon$ and $\alpha^{2}$ on $\kappa$ for these new solutions are opposite to those for $\omega=$0. Since the $\tau$-growth increases a spectral loss, it causes the $\alpha^{2}$-decrease and $\Upsilon$-rise (compare Figs. 3 and 4).
@@
Figure 2: Dimensionless pulse width $\Upsilon$ (a) and intensity $\alpha_{0}^{2}$ (b) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=$-0.01, $\tau=$0.01. The unstable solutions are shown by dashed curves. A bold magenta curve shows the stable solutions for $\omega=10^{-5}$.
Figure 3: Dimensionless pulse width $\Upsilon$ (a) and intensity $\alpha_{0}^{2}$ b) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=-$0.001, $\tau=$0.01. Dashed curves correspond to unstable solutions. Solutions corresponding to $\omega$=10-4 and 10-3 are nonphysical.
Figure 4: Dimensionless pulse width $\Upsilon$ (a) and intensity $\alpha_{0}^{2}$ (b) in dependence on the transverse ($\kappa$) and longitudinal ($\omega$) gain grading parameters. $\Lambda^{\prime}=-$0.001, $\tau=$0.1. Insets show the analytical spatio-temporal (a) and spectral (b) profiles of STDS for the parameters marked by the green circles ($\kappa=$0.0006, $\omega=$10-7). The blue square marks the parameters $\kappa=$0.0005, $\omega=$10-6 corresponding to Fig. 7. The unstable solutions are shown by dashed curves. The white square marks the parameters corresponding to Fig. 8. Solutions for $\omega$=10-4 and 10-3 are nonphysical.
2.4 STDS stability: linear analysis
The linear stability analysis is based on the standard technique of the study of the Jacobian eigenvalues (see Appendix C and Kalashnikov (2024)) at the stationary points of (13). Their real parts must be negative to provide a local stability of the STDS solutions.
The analysis demonstrates that the solutions presented in Fig. 2 are unstable, expecting the confined region of $\kappa\to$0 for a non-zero $\omega$ (thick magenta line). The dependencies of the corresponding Jacobian eigenvalues $\lambda$ on $\kappa$ are shown in Fig. 5.
When $\Lambda^{\prime}\to$0, the additional unstable solutions appear for the considered non-zero $\omega$ as is shown in Figs. 3,4. But in this case, the first solutions corresponding to the branches with lower $\alpha^{2}$ are stable for $\omega\neq$0 within all diapason of $\kappa$ in which an STDS exists. This diapason shrinks with $\omega$ because the $\omega$-growth shortens a gain “window” in the $t^{\prime}$-domain.
The analysis testifies that, for an unstable solution, there is always an eigenvalue $\lambda$ with $\Re(\lambda)>$0 but $\Im(\lambda)=$0 that does not show evidence of the limiting circles. On the other side, the numerical simulation of (13) demonstrates a rapid convergence to a stable solution within its existence region, as is shown in Fig. 6. One must note that an absence of analytical solutions does not mean a lack of localized solutions because the model is confined by the choice of the ansatz (7). Therefore, the numerical simulations of Eq. (2) are required.
Figure 5: Real and imaginary parts of the Jacobian eigenvalues $\lambda$ in dependence on $\kappa$ for $\omega$=10-5, $\Lambda^{\prime}=-$0.01, and $\tau=$0.01 Kalashnikov (2024).
Figure 6: Evolution of $\Upsilon(Z)$ and $\alpha^{2}(z)$ for the parameters marked by points in Fig. 4. The initial values are $\Upsilon(0)=$10, $\alpha(0)=$0.1, $\rho(0)=$1, and $\psi(0)=\theta(0)=\phi(0)=$0.
2.5 STDS stability: numerical simulations
@@
We performed large-scale numerical simulations of Eq. (2) based on the COMSOL Multiphysics software to analyze the stability of the analytical solutions based on the VA. The characteristic case is illustrated in Fig. 7 for the parameters marked by the blue square in Fig. 4. The red line corresponds to the analytical solution, and the black curve shows the numerical solution for the maximal $|\psi|^{2}$ in dependence on $Z$. The ansatz (7) with $\alpha_{0}$=0.46, $\Upsilon$=22, $\rho$=1, $\psi=\theta$=0 was chosen as the initial condition. One can see a fast approach of the STDS maximal intensity to the analytical solution.
Figure 7: Evolution of the STDS peak intensity $|\psi(R=0,t^{\prime}=0,Z)|^{2}$ (black solid curve) and the STDS profile $|\psi(R,t^{\prime},Z=3000)|^{2}$ for the parameters of Fig. 4 (blue square). The red line shows the analytical $\alpha_{0}^{2}$ from Fig. 4 for $\omega=$10-6, $\kappa=$0.0005.
Besides the stable STDS solutions, the numerical simulations demonstrate two main destabilization scenarios. The first is the STDS decay (Fig. 8) corresponding to an unstable analytical solution of Fig. 4,b (white square). For this scenario, a peak amplitude decreases exponentially with a slow temporal broadening of STDS.
Figure 8: Evolution of the STDS peak intensity $|\psi(R=0,t^{\prime}=0,Z)|^{2}$ (black solid curve, logarithmic scale), and the initial and final STDS profiles $|\psi(R,t^{\prime},Z=0,10000)|^{2}$ for the parameters of Fig. 4 marked by the white square.
@@
The second scenario is the STDS splitting with the irregular oscillations between multiple pulse states leading to a “rogue wave” dynamics. Fig. 9 shows such a regime, which cannot be described within the above analytical model. One may conjecture that the splitting in a local time domain could be explained by broadening the confinement potential with the $|\Lambda^{\prime}|$-growth. However, a more precise analysis of such dynamics requires considering the gain saturation.
Figure 9: Evolution of the maximal $|\psi|^{2}_{max}$ with $Z$ for $\omega=10^{-5}$, $\kappa=$0.005, $\Lambda^{\prime}=$-0.01, and $\tau=$0.1. Inset shows the profile of the STDS complex at the final point of the considered $Z$-interval.
@@ 2.6 STDS stability under the saturable gain dynamics
To take into account the dynamic gain saturation, it is convenient to use the Frantz-Nodvik equation Frantz and Nodvik (1963):
@@
where $G_{ss}$ is a small signal gain, $T_{r}$ is a gain relaxation time, $E_{s}=h\nu/S\sigma_{em}$ is a gain saturation energy ($\nu$ is a central gain-band wavelength coinciding with the STDS carrier frequency, $S$ is a beam area, $\sigma_{em}$ is a stimulated emission cross-section). Specifically, we base on the characteristics of a Cr:ZnS waveguide laser Sorokin et al. (2022). Table 3 presents the corresponding dimensional laser parameters ($n_{2}$ value is defined at 1.3 $\mu$m).
Table 3: Cr:ZnS laser parameters Sorokina et al. (2002); Sorokina (2008)
parameters values
radiative lifetime $T_{r}$, $\mu s$ 4.3
@@
central wavelength $\lambda$, $\mu$m 2.35
@@
$n_{0}$ at 2.3 $\mu$m 2.27
$\delta$ 0.0047
numerical aperture $NA=\sqrt{n_{1}^{2}-n_{0}^{2}}$ $\approx 0.15$
@@
For a chosen level of linear loss $\Lambda$, some characteristic profiles of STDS based on the numerical simulations are shown in Figs. 10. As one can see, an additional scenario of STDS destabilization appears. Namely, there are the STDS shape distortion, shift, and splitting, which alternate irregularly.
Figure 10: The $|\psi|^{2}$-profiles on the different distances $Z=$16000 (a), 29000 (b), and 30000 (c) for $\kappa=$0.002, $\omega=10^{-5}$, $\Lambda$=0.1, $G_{ss}$=1.1 $\Lambda$, and $\tau=$0.1 (see the green circle in Fig. 11).
Figure 11: The visualization of the STDS stability region on ($\kappa$$\omega$)-plane under dynamic gain saturation. The nodes correspond to the “last” numerical stability points connected “by hand” for clarity. The green circle corresponds to the parameters of Fig. 10. $\Lambda$=0.1, $G_{ss}$=1.1 $\Lambda$, and $\tau=$0.1.
The region of STDS stability is shown in Fig. 11 for a given linear loss coefficient $\Lambda$. The “nodes” in Figure correspond to the “last” stability points obtained numerically and are connected “by hand” for clarity. The STDS decay confines this region for larger $\kappa$, so there are no stable pulses for $\kappa\gtrsim 0.006$. The difference with the analytical results lies in the reduction of a net gain required for the soliton formation. Such a gain reduction is a direct consequence of the gain saturation. The main destabilization factor for smaller $\kappa$ is irregular pulse splitting, which alternates with a single-pulse regime (Figs. 10). The asymmetry of the STDS complex appears due to dynamic gain saturation (Figs. 10). When $\kappa$ approaches zero, very strong STDS power and width oscillations appear. They reach more than two orders of magnitude (Figs. 10 (a) vs. (b,c)), so STDS is periodically “reborn”. This phenomenon confines the stability region from below on $\kappa$ and above on $\omega$.
As an additional effect that waits for further exploration is a symmetry breaking. It was found Kalashnikov and Wabnitz (2022) that this effect contributes to STDS dynamics in a system with pure real pancake-like plus imaginary cigar-like potentials. The STDS shape can be distorted, and the dynamics becomes more complicated. However, STDS remains 3D-localized. The situation is reversed in our case: pure imaginary pancake-like plus real cigar-like potentials. The numerical analysis shows that, unlike Kalashnikov and Wabnitz (2022), there is no STDS splitting in the transverse dimension $R$. That is, the transverse perturbations are suppressed, and we may conjecture that the development of the spatial asymmetry would be troubled.